OpenStudy (anonymous):

Find the slope of the line that passes through the points (5, 8) and (9, 4). A. 3 B. 1 C. -1 D. four-thirds

OpenStudy (anonymous):

(4-8)/(9-5) = -4/4 = -1

OpenStudy (hoblos):

C